# Learning Noise-Robust Stable Koopman Operator for Control with Hankel DMD

### v5

@@ I Introduction
Recently, the concept of polyflow was introduced by [25]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept is closely related to the derivative-based approximation of Koopman operator [26] except closure is considered. In earlier work, it was observed by Mezić [27] that instead of relying on arbitrary sets of observables, one could use observables generated by the dynamics, such as time delays that form a Krylov subspace, to study the spectral properties of the Koopman operator. This approach has also been applied in the context of Hankel-DMD methods. Despite the initial success of Polyflow for MPC [28], merely minimizing the one-step-ahead error remains susceptible to noise in the data. The second challenge is to ensure long-term prediction accuracy. Wu et al. [29] revealed that even under mild noise and nonlinearity, DMD struggles to accurately predict system dynamics. Several recent approaches have been proposed to address the challenges posed by measurement noise in learning Koopman operators. Sinha et al. [30] developed an online robust Koopman framework that employs a min-max optimization strategy to mitigate the effect of noisy streaming data, while Hao et al. [31] proposed a deep Koopman learning method that incorporates bounded noise considerations by updating the observable functions during training. However, these methods either require restrictive assumptions on noise bounds or face challenges in ensuring long-term prediction stability, which motivates the need for our stable parameterization and rollout-based learning strategy. Although the use of roll-out techniques for recurrent loss has proven beneficial when data are limited and noisy both in the deep learning [32, 33, 34] and the reduced-order modeling community [35], optimizing recurrent loss can be tricky and prone to poor local minimum in practice. The third challenge in Koopman operator-based control is to ensure the stability of the Koopman operator itself, since the real part of any continuous-time eigenvalue of the Koopman operator of any physical system should not be positive [34, 36]. If the forward pass becomes unstable, it inevitably destabilizes the backward pass as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling especially in the high-dimensional phase space and, consequently, poor control performance. Besides, it is important to note that we do not attempt to address other challenges such as the convergence of Koopman operator. Interested readers are referred to this excellent paper by Colbrook and Mezić [37].
@@
In order to address the above issues, the goal of this paper is to develop a noise-robust learning framework for the Koopman operator. Specifically, the contributions of this work are:
@@ II Proposed Framework
II-A Lifted Linear Predictor
Consider a nonlinear discrete-time dynamical system,
$x_{k+1}=f(x_{k},u_{k}),$ (1)
where $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ represents a general nonlinear mapping, and $x_{k}=x(t_{k})\in\mathbb{R}^{n}$ and $u_{k}=u(t_{k})\in\mathbb{R}^{m}$ denote the system state and control input at time $t_{k}$, respectively. Given an initial condition $x_{0}$ at time $t_{0}$ and a sequence of control inputs $\{u_{k}\}_{k=0}^{N_{p}}$, where $N_{p}$ is the length of the prediction horizon, the objective is to estimate the sequence of future states $\{x_{k}\}_{k=1}^{N_{p}+1}$.
$z_{0}$ $=\Phi(x_{0}),$ (2)
$z_{k+1}$ $=Az_{k}+Bu_{k},$ $k=0,1,\dots,N_{p},$
$\hat{x}_{k}$ $=Cz_{k},$ $k=1,\dots,N_{p}+1.$
where $z_{k}\in\mathbb{R}^{N_{\phi}}$ is the lifted state in an extended space of dimension $N_{\phi}$, and $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{N_{\phi}}$ is a nonlinear lifting function that maps the original state space to a higher-dimensional feature space. The vector $\hat{x}_{k}\in\mathbb{R}^{n}$ represents the estimated state of the system at time $t_{k}$.
The lifting function $\Phi(\cdot)$ is defined as:
where each function $\phi_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is an observable, capturing potentially nonlinear transformations of the state. The choice of observables $\{\phi_{i}(x)\}_{i=1}^{N_{\phi}}$ determines the accuracy of the linear approximation.
It is important to note that, under certain conditions (e.g., control-affine systems), one can also adopt a bilinear model realization [38, 39, 40, 41, 42]. Suppose there exist coefficient matrices $A\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad B_{i}\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad C\in\mathbb{R}^{n\times N_{\phi}},$ such that for $z_{k}=[z_{k}^{(1)},\dots,z_{k}^{(N_{\phi})}]^{\top},\quad x_{k}=[x_{k}^{(1)},\dots,x_{k}^{(n)}]^{\top},\quad u_{k}=[u_{k}^{(1)},\dots,u_{k}^{(m)}]^{\top},$ the dynamics takes the form
$z_{k+1}$ $=Az_{k}+\sum_{i=1}^{m}u_{k}^{(i)}B_{i}z_{k},$ (4)
$x_{k}$ $=C\,z_{k},$
then $\bigl{(}A,\{B_{i}\}_{i=1}^{m},C\bigr{)}$ is said to be a bilinear realization of the system. While extending our approach to the bilinear setting is natural, we opt here for a linear formulation for simplicity and seamless integration with standard control methods.
II-B Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor via state-space lifting , we briefly outline the Koopman operator approach for analyzing a controlled dynamical system, following the approach of [15]. We consider the following discrete-time nonlinear system with control,
@@ II-B Koopman Operator for Controlled System
where, $x\in\mathbb{R}^{n}$ is the state of the system and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the state of the system for the next time step. We define the extended state comprising of the state $x$ of system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$. Despite the fact that the Koopman operator $\mathcal{K}$ is a linear operator that fully describes the time evolution of any observable function in a non-linear dynamical system, it is inherently infinite-dimensional which makes learning the Koopman operator challenging.
II-C Extended DMD for Control with Dynamics-Driven Observables
@@ II-D1 Known Dynamics
When the system dynamics $f(x,u)$ is explicitly known, we define the Polyflow basis functions of order $N$, denoted as $T_{N}(x)$, as follows:
@@
Here, $f^{(i)}(x,0)$ represents the $i$-th recursive application of $f$ with zero input, capturing the intrinsic system evolution. The dimension of the lifted state is $N_{\phi}=N\times n$.
Using these Polyflow basis functions to lift the system state, we obtain the following linear system representation, consistent with the lifted linear predictor in eq. 2:
@@
Connection to Time-Delay Embedding
While both time-delay embedding [44] and Polyflow lift the state into a higher-dimensional space, their methodologies differ. Time-delay embedding constructs an extended state vector by recording historical state observations over multiple timesteps, effectively capturing past system behavior. In contrast, Polyflow directly leverages the governing equations to generate an instantaneous lifted representation without requiring historical state trajectories. This fundamental distinction makes Polyflow particularly suitable for embedding systems where access to past observations is limited or infeasible.
@@ II-E Roll-out Loss
To improve long-term prediction accuracy and noise resistance in training data, we employ a multi-step rollout loss instead of the one-step prediction loss eq. 8.
Consider a data set consisting of $M$ trajectories, each trajectories consisting of $R$ roll-out steps, $\{(x_{0,j},x_{1,j},\dots,x_{R,j},u_{0,j},\dots,u_{R-1,j})\}_{j=1}^{M}$, where $x_{i,j}$ denotes the state of the system at the $i$th time step of the $j$-th trajectory, and $u_{i,j}$ is the corresponding control input.
The optimization problem for learning the system matrices $A$ and $B$ is then formulated as:
@@
where $A$ and $B$ are the unknown system matrices to be learned, and $T_{N}(x)$ represents the Polyflow basis function transformation of the state.
The hyperparameters $N$, and $R$ are typically determined through hyperparameter tuning to balance model complexity and predictive performance. Adjusting these parameters is a trade-off: increasing $N$ can capture more complex dynamics but may lead to overfitting if not managed carefully, while higher $R$ values typically enhance long-term prediction accuracy at the cost of increased computational burden.
@@ II-F Curriculum Learning
In our framework, we implement a form of curriculum learning by progressively increasing the length of the rollout $R$ during the training process. Initially, the model is trained with shorter rollout lengths, focusing on immediate predictions. As training progresses and the model stabilizes, the roll-out length $R$ increases geometrically after certain epochs. This gradual increase is expected to allow the model to learn from easier short-term predictions before tackling the more complex task of long-term prediction, thus aligning with the curriculum learning paradigm. In doing so, the model is expected to build its capacity to handle extended sequences, improving overall prediction accuracy over long time horizons. Compared to direct exposure to long rollout lengths from the start, this progressive approach offers optimization benefits by improving stability and convergence. Training with long roll-outs early on can lead to high variance in gradients and poor local optima due to compounding prediction errors. By starting with shorter roll-outs, the model establishes a more robust foundation before incrementally handling extended sequences, ultimately enhancing prediction accuracy over long time horizons.
@@ II-G Stabilized Formulation
To guarantee the stability of both the Koopman matrix $A$ and the training process under rollout loss, we enforce stable eigenvalues of $A$ employing the fact that a skew-symmetric matrix with negative real-valued diagonal elements has stable or neutral stable eigenvalues in continuous-time dynamical system [34]. This guarantees the stability of the Koopman matrix during training with roll-out loss, preventing instability in both forward and backward passes.
@@
$\hat{A}=\exp\left(\tilde{A}\Delta t\right).$ (14)
Since the Koopman operator is typically expressed in a different coordinate system, a similarity transformation using a coordinate transformation matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is required to recover $A$:
@@ II-H Model Predictive Control
One of the most important applications of lifted linear predictor is to provide a surrogate model for MPC, which is referred as Koopman-MPC (KMPC) [18]. At each discrete time step, the MPC solves the following optimization problem over the prediction horizon $N_{p}$:
@@
II-I Iterative Learning
Models trained on trajectories that are generated by randomly sampling the phase space and control input could perform poorly when subjected to optimal feedback control input. This is mainly because these trained models lack exposure to the desired scenarios during the training process, especially for a relatively high-dimensional system. To address this issue, we propose a straightforward iterative data augmentation technique inspired by the work of Uchida and Duraisamy [46]. The idea is to iteratively collect the response of the true system under control input synthesized with the previous data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the data-driven model. Our rationale is that such experience from the true system can be used to better explore the phase space of the nonlinear system than merely ad hoc random sampling of the phase space, which is subject to the curse of dimensionality. Hence, our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator). A summary of the proposed framework is illustrated in the figure 1.
Figure 1: Overview of the roll-out EDMD Koopman Model. The process begins with trajectory data collected from a nonlinear dynamical system. The states are lifted through Polyflow basis functions to obtain a higher-dimensional representation. This representation is then used in the roll-out EDMD algorithm to derive a linear model. The resulting linear model is employed for Model Predictive Control (MPC), which is iteratively refined through iterative learning from closed loop trajectories.

### v6

@@ I Introduction
Recently, the concept of polyflow was introduced by [25]. Polyflow employs the governing equations of the system to construct basis functions, offering a non-local linearization approach for nonlinear differential equations. This concept is closely related to the derivative-based approximation of Koopman operator [26] except closure is considered. In earlier work, it was observed by Mezić [27] that instead of relying on arbitrary sets of observables, one could use observables generated by the dynamics, such as time delays that form a Krylov subspace, to study the spectral properties of the Koopman operator. This approach has also been applied in the context of Hankel-DMD methods. Despite the initial success of Polyflow for MPC [28], merely minimizing the one-step-ahead error remains susceptible to noise in the data. The second challenge is to ensure long-term prediction accuracy. Wu et al. [29] revealed that even under mild noise and nonlinearity, DMD struggles to accurately predict system dynamics. Several recent approaches have been proposed to address the challenges posed by measurement noise in learning Koopman operators. Sinha et al. [30] developed a robust online Koopman framework that employs a min-max optimization strategy to mitigate the effect of noisy streaming data, while Hao et al. [31] proposed a deep Koopman learning method that incorporates considerations of bounded noise by updating observable functions during training. However, these methods either require restrictive assumptions on noise bounds or face challenges in ensuring long-term prediction stability, which motivates the need for our stable parameterization and rollout-based learning strategy. Although the use of roll-out techniques for recurrent loss has proven beneficial when data are limited and noisy both in the deep learning for the Koopman operator [32, 33, 34] and the reduced-order modeling community [35], optimizing recurrent loss can be tricky and prone to poor local minimum in practice. The third challenge in Koopman operator-based control is to ensure the stability of the Koopman operator itself, since the real part of any continuous-time eigenvalue of the Koopman operator of any physical system should not be positive [34, 36, 37]. If the forward pass becomes unstable, it inevitably destabilizes the backward pass as well. Lastly, curating training data for Koopman operator-based control often depends on ad-hoc selections of input signals and initial conditions, which can result in insufficient sampling especially in the high-dimensional phase space and, consequently, poor control performance. Besides, it is important to note that we do not attempt to address other challenges such as the convergence of Koopman operator. Interested readers are referred to this excellent paper by Colbrook and Mezić [38].
@@
In order to address the above issues, the goal of this paper is to develop a noise-robust learning framework for the Koopman operator. Specifically, the contributions of this work are as follows:
@@ II Proposed Framework
II-A Koopman Operator for Controlled System
To provide a rigorous justification for constructing a linear predictor by lifting the state space, we briefly outline the Koopman operator approach to analyze a controlled dynamical system, following the approach of [15]. We consider the following discrete-time nonlinear system with control,
@@ II-A Koopman Operator for Controlled System
where $x\in\mathbb{R}^{n}$ is the state of the system, $x^{+}$ is the system state at the next time step, and $u\in\mathcal{U}\subset\mathbb{R}^{m}$ is the control input. $\mathcal{U}$ is the space of the available control inputs. $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\rightarrow\mathbb{R}^{n}$ is the transition mapping. $x^{+}$ is the state of the system for the next time step. We define the extended state comprising the state $x$ of the system and the control sequence $\mathbf{u}\in\ell(\mathcal{U})$, where $\ell(\mathcal{U})$ is the space of all sequences $(u_{i})_{i=0}^{\infty}$ with $(u_{i})_{i=0}^{\infty}\in\mathcal{U}$,
@@
for any observable $\psi:\mathbb{R}^{n}\times\ell(\mathcal{U})\to\mathbb{R}$ belonging to some space of observables $\mathcal{H}$. Despite the fact that the Koopman operator $\mathcal{K}$ is a linear operator that fully describes the time evolution of any observable function in a non-linear dynamical system, it is inherently infinite-dimensional which makes learning the Koopman operator challenging. In the following subsection, we adopt the widely used concept of lifted linear predictor [39, 40] as a finite-dimensional approximation of the Koopman operator.
II-B Lifted Linear Predictor
Let us use a subscript to indicate the time step and rewrite eq. 1 as follows.
$x_{k+1}=f(x_{k},u_{k}),$ (3)
where $f:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}^{n}$ represents a general non-linear mapping, and $x_{k}=x(t_{k})\in\mathbb{R}^{n}$ and $u_{k}=u(t_{k})\in\mathbb{R}^{m}$ denote the state of the system and the input of the control at time $t_{k}$, respectively. Given an initial condition $x_{0}$ at time $t_{0}$ and a sequence of control inputs $\{u_{k}\}_{k=0}^{N_{p}}$, where $N_{p}$ is the length of the prediction horizon, the objective is to estimate the sequence of future states $\{x_{k}\}_{k=1}^{N_{p}+1}$.
$z_{0}$ $=\Phi(x_{0}),$ (4)
$z_{k+1}$ $=Az_{k}+Bu_{k},$ $k=0,1,\dots,N_{p},$
$\hat{x}_{k}$ $=Cz_{k},$ $k=1,\dots,N_{p}+1.$
where $z_{k}\in\mathbb{R}^{N_{\phi}}$ is the lifted state in an extended space of dimension $N_{\phi}$, and $\Phi:\mathbb{R}^{n}\to\mathbb{R}^{N_{\phi}}$ is a non-linear lifting function that maps the original state space to a higher dimensional feature space. The vector $\hat{x}_{k}\in\mathbb{R}^{n}$ represents the estimated state of the system at time $t_{k}$.
The lifting function $\Phi(\cdot)$ is defined as:
where each function $\phi_{i}:\mathbb{R}^{n}\to\mathbb{R}$ is an observable, capturing potentially non-linear transformations of the state. The choice of observables $\{\phi_{i}(x)\}_{i=1}^{N_{\phi}}$ determines the accuracy of the linear approximation.
It is important to note that, under certain conditions (e.g., control-affine systems), one can also adopt a bilinear model realization [41, 42, 43, 44, 45]. Suppose that there exist coefficient matrices $A\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad B_{i}\in\mathbb{R}^{N_{\phi}\times N_{\phi}},\quad C\in\mathbb{R}^{n\times N_{\phi}},$ such that for $z_{k}=[z_{k}^{(1)},\dots,z_{k}^{(N_{\phi})}]^{\top},\quad x_{k}=[x_{k}^{(1)},\dots,x_{k}^{(n)}]^{\top},\quad u_{k}=[u_{k}^{(1)},\dots,u_{k}^{(m)}]^{\top},$ the dynamics takes the form
$z_{k+1}$ $=Az_{k}+\sum_{i=1}^{m}u_{k}^{(i)}B_{i}z_{k},$ (6)
$x_{k}$ $=C\,z_{k},$
then $\bigl{(}A,\{B_{i}\}_{i=1}^{m},C\bigr{)}$ is said to be a bilinear realization of the system. Although extending our approach to the bilinear setting is natural, here we opt for a linear formulation in eq. 4 for simplicity and seamless integration with standard control methods [46].
II-C Extended DMD for Control
@@ II-D1 Known Dynamics
When the dynamics of the system $f(x,u)$ is explicitly known, we define the Polyflow basis functions of order $N$, denoted as $T_{N}(x)$, as follows:
@@
Here, $f^{(i)}(x,0)$ represents the $i$-th recursive application of $f$ with zero input, which captures the intrinsic evolution of the system. The dimension of the lifted state is $N_{\phi}=N\times n$.
Using these Polyflow basis functions to lift the state of the system, we obtain the following linear system representation, consistent with the linear predictor lifted in eq. 4:
@@ II-D2 Unknown Dynamics
II-D3 Connection to Hankel DMD
Our proposed use of Polyflow basis functions bears conceptual similarity to the Hankel DMD method [48, 49], as both leverage delay-embedded observables to approximate the action of the Koopman operator through a Krylov subspace structure. Specifically, the Polyflow recursion we employ mimics delay embeddings by iteratively composing either the true or learned dynamics to generate basis functions. However, unlike Hankel DMD [49]—which constructs purely numerical delay-coordinate snapshots from raw trajectory data—Polyflow builds its function basis through recursive application of (possibly learned) dynamics, resulting in a physics-informed lifting map in an EDMD fashion. This makes Polyflow particularly advantageous when historical data is unavailable for starting.
@@ II-E Roll-out Loss
To improve long-term prediction accuracy and noise resistance in training data, we employ a multi-step roll-out loss instead of the one-step prediction loss eq. 8. The multi-step roll-out loss refers to cumulative prediction error in multiple future time steps (also known as rollout length). The rollout length $R$ determines how many steps in the future the model is trained to predict.
Consider a data set consisting of $M$ trajectories, each trajectories consisting of $R$ roll-out steps, $\{(x_{0,j},x_{1,j},\dots,x_{R,j},u_{0,j},\dots,u_{R-1,j})\}_{j=1}^{M}$, where $x_{i,j}$ denotes the state of the system at the $i$ -th time step of the $j$-th trajectory, and $u_{i,j}$ is the corresponding control input.
The optimization problem for learning the system matrices $A$ and $B$ is then formulated as follows:
@@
where $A$ and $B$ are the unknown system matrices to be learned, and $T_{N}(x)$ represents the transformation of the Polyflow basis function of the state of the system.
The hyperparameters $N$ and $R$ are typically determined by tuning the hyperparameters to balance the complexity of the model and predictive performance. Adjusting these parameters is a trade-off: increasing $N$ can capture more complex dynamics but can lead to overfitting if not managed carefully, while higher values of $R$ typically enhance long-term prediction accuracy at the cost of increased computational burden.
@@ II-F Curriculum Learning
In our framework, we implement a form of curriculum learning by progressively increasing the length of the rollout $R$ during the training process. Progressive scheduling refers to gradually increasing the complexity of the training task - in our case, by increasing the length of roll-out $R$ during training. This aligns with the principles of curriculum learning, where the model first learns easier subtasks before tackling harder ones. We refer to this approach as progressive rollout schedule.
Initially, the model is trained with shorter rollout lengths, focusing on immediate predictions. As training progresses and the model stabilizes, the roll-out length $R$ increases geometrically after certain epochs. This gradual increase is expected to allow the model to learn from easier short-term predictions before tackling the more complex task of long-term prediction, thus aligning with the curriculum learning paradigm. In doing so, the model is expected to build its capacity to handle extended sequences, improving overall prediction accuracy over long time horizons. Compared to direct exposure to long rollout lengths from the start, this progressive approach offers optimization benefits by improving stability and convergence. Training with long roll-outs early on can lead to high variance in gradients and poor local optima due to compounding prediction errors. By starting with shorter roll-outs, the model establishes a more robust foundation before incrementally handling extended sequences, ultimately enhancing prediction accuracy over long time horizons.
@@ II-G Stabilized Formulation
To guarantee the stability of both the Koopman matrix $A$ and the training process under rollout loss, we enforce the eigenvalues of $A$ to be stable, employing the fact that a skew-symmetric matrix with negative real-valued diagonal elements has stable or neutral stable eigenvalues in continuous-time dynamical system [34]. This guarantees the stability of the Koopman matrix during training with roll-out loss, preventing instability in both forward and backward passes.
@@
$\hat{A}=\textrm{expm}\left(\tilde{A}\Delta t\right),$ (14)
where expm refers to matrix exponential. Since the Koopman operator is typically expressed in a different coordinate system, a similarity transformation using a coordinate transformation matrix $P\in\mathbb{R}^{\tilde{n}\times\tilde{n}}$ is required to recover $A$:
@@ II-H Model Predictive Control
One of the most important applications of the lifted linear predictor is to provide a surrogate model for MPC, which is known as Koopman-MPC (KMPC) [18]. At each discrete time step, the MPC solves the following optimization problem on the prediction horizon $N_{p}$:
@@
II-I Iterative Data Augmentation
Models trained on trajectories that are generated by randomly sampling the phase space and control input could perform poorly when subjected to optimal feedback control input. This is mainly because these trained models lack exposure to the desired scenarios during the training process, especially for a relatively high-dimensional system. To address this issue, we propose a straightforward iterative data augmentation technique inspired by the work of Uchida and Duraisamy [51]. The idea is to iteratively collect the response of the true system under control input synthesized with the previous data-driven model. Such response together with control input is used to augment the original dataset to further refine and improve the data-driven model. Our rationale is that such an experience from the true system can be used to better explore the phase space of the nonlinear system than merely ad hoc random sampling of the phase space, which is subject to the curse of dimensionality.
Hence, our proposed framework is named as STIRK (Stable Training with Iterative Roll-out Koopman Operator) as illustrated in figure fig. 1.
Figure 1: Overview of the Stable Training with Iterative Roll-out Koopman Operator framework. The process begins with trajectory data collected from a nonlinear dynamical system. The states are lifted through Polyflow basis functions in eq. 9 if the system dynamics is known, or a recursive structure in eq. 11 if the system dynamics is unknown, to obtain a higher-dimensional representation. This representation is then used in the roll-out EDMD algorithm to derive a linear model. The resulting linear model is employed for Model Predictive Control (MPC), which is further retrained through iterative data augmentation from closed loop trajectories.